OpenStudy (anonymous):

find the midpoint between p(-4)and Q(8) on a number line.

OpenStudy (anonymous):

pleasee help meee

OpenStudy (anonymous):

(-4 +8)/2 =4/2 =2

OpenStudy (anonymous):

then you can make sure it makes sense (going from -4 to 2 on the numberline is 6 steps. going from 2 to 8 on the numberline is also 6 steps, so this is a midpoint! it's the middle!